OpenStudy (anonymous):

Write the nth term of the following sequence in terms of the first term of the sequence. 2, -4, 8, -16, . . .

OpenStudy (amistre64):

ignore the alternating for the moent; is there a pattern you can pick up on?

OpenStudy (anonymous):

Multiplying by 2

OpenStudy (amistre64):

good, then we need to construct a recurrsion based on that knowledge i believe

OpenStudy (amistre64):

\[A_{n+1}=2*A_n\] but to include the alternating negative lets make it a -2\[A_{n+1}=-2*A_n\]

OpenStudy (amistre64):

using this we can work it backwards to the A0 term

OpenStudy (amistre64):

or we can go forwards with what is known .....

OpenStudy (amistre64):

a1 = 2 a2 = a1*-2 a3 = a1*-2*-2 = a1 * (-2)^2 a4 = a1*(-2^2)*(-2) = a1*(-2)^3 if i assume a pattern a5 = a1 * (-2)^4 a6 = a1 * (-2)^5 an = a1 * (-2)^(n-1)

OpenStudy (anonymous):

Is the answer -(-2)^n?

OpenStudy (amistre64):

"in terms of the first term of the sequence" whatever that my entail but perhaps if we simplified it

OpenStudy (anonymous):

how do you do it in terms of the first sequence?

OpenStudy (amistre64):

replace a1 by its value of 2 im not sure if simplifying is what they want or not tho ....

OpenStudy (anonymous):

do you know a trick to turn the sequence for all terms into a term for the first sequence?

OpenStudy (anonymous):

What? what a1?

OpenStudy (amistre64):

i do, there is a backwards version and a forwards version; i just showed you the forward version

OpenStudy (amistre64):

a1 is the first term of the seq ..

OpenStudy (amistre64):

an = a1 * (-2)^(n-1) ^^

OpenStudy (anonymous):

I don't understand this at all. can you help me with a, a + 4, a + 8, . . . That's 10 times the hell this question is

OpenStudy (anonymous):

I don't even understand what the answer is

OpenStudy (anonymous):

:"(

OpenStudy (amistre64):

swearing at me doesnt make me want to help you much :/

OpenStudy (anonymous):

i'm not swearing at you. Sorry, i'm very frustrated.

OpenStudy (amistre64):

if we separate the terms term : value 1: a 2: a+4 3: a+4+4 4: a+4+4+4 5: a+4+4+4+4 etc ....

OpenStudy (anonymous):

I'll go relearn this instead. Thanks for your help.

OpenStudy (amistre64):

we can then simplfy the results term : value 1: a + 0(4) 2: a + 1(4) 3: a + 2(4) 4: a + 3(4) 5: a + 4(4) notice that each term follows a pattern such that n: a + (n-1)(4)

OpenStudy (amistre64):

without knowing how the grading goes, thats prolly the best i can assess it